# thanks: Corresponding author

### v1

@@ acknowledgment
Appendix A Proof of Eq. (49)
Vectors given by Eq. (46) are linearly independent and span an L(d1)L(d-1)-dimensional subspace of $\mathcal{H}\otimes\mathcal{H}$. Accordingly, corresponding to the nonorthogonal vectors (46), one can define the reciprocal vectors $\{|\widetilde{\phi}_{k}^{(\alpha)}\rangle\}$ such that $\langle\phi_{k}^{(\alpha)}|\widetilde{\phi}_{k^{\prime}}^{(\alpha^{\prime})}\rangle=\delta_{kk^{\prime}}\delta^{(\alpha\alpha^{\prime})}$. The reciprocal vectors can be expanded in terms of the original vectors $|\phi_{k}^{(\alpha)}\rangle$ as
$=$ $\left[CG^{\textrm{T}}\right]_{kl}^{(\alpha\beta)}.$
This implies that $CG^{\textrm{T}}={\mathbb{I}}_{L(d-1)}$, or equivalently, $C^{\textrm{T}}=G^{-1}$, where T stands for the transposition. Accordingly, up to a transposition, the matrix of the coefficients is obtained by the inverse of the Gram matrix $G$. More precisely
$|\widetilde{\phi}_{k}^{(\alpha)}\rangle=\sum_{k^{\prime}=1}^{d-1}\sum_{\alpha^{\prime}=1}^{L}[G^{-1}]_{k^{\prime}k}^{(\alpha^{\prime}\alpha)}|\phi_{k^{\prime}}^{(\alpha^{\prime})}\rangle.$ (67)
Now, for a given qudit quantum state $\varrho$, we define a two-qudit pure state $\varrho\otimes{\mathbb{I}}]|\phi\rangle$, and expand it in terms of the aforementioned basis of $\mathcal{H}\otimes\mathcal{H}$, we get
$[\varrho\otimes{\mathbb{I}}]|\phi\rangle=\frac{1}{d}|\phi\rangle+\sum_{\alpha=1}^{L}\sum_{k=1}^{d-1}a_{k}^{(\alpha)}|\phi_{k}^{(\alpha)}\rangle+|\chi\rangle,$ (68)
where the first term follows from $\langle\phi|[\varrho\otimes{\mathbb{I}}]|\phi\rangle=1/d$, in conjunction with Eq. (45). The unnormalized vector $|\chi\rangle$ represents the component of $\varrho\otimes{\mathbb{I}}]|\phi\rangle$ that lies in the orthogonal subspace spanned by $\{|\chi_{l}\rangle\}$. For the second term, we have
$a_{k}^{(\alpha)}$ $=$ $\langle\widetilde{\phi}_{k}^{(\alpha)}|[\varrho\otimes{\mathbb{I}}]|\phi\rangle$ (69)
$=$ $\frac{1}{d}\sum_{\beta=1}^{L}\sum_{l=1}^{d-1}[G^{-1}]_{kl}^{(\alpha\beta)}\sum_{j=0}^{d-1}{\omega^{\ast}}^{lj}p_{j}^{(\beta)}$
$=$ $[G^{-1}\Omega]_{k}^{(\alpha)},$
[…]
Appendix B Derivation of the eigenvalues of $G$
@@
$D_{L}$ $=$ ${\mathrm{diag}}\{L-1,-1,\cdots,-1\},$ (73)
$\Lambda_{d-1}$ $=$ ${\mathrm{diag}}\{\lambda_{1},-\lambda_{1},\cdots,\lambda_{\lfloor\frac{d-1}{2}\rfloor},-\lambda_{\lfloor\frac{d-1}{2}\rfloor},\lambda_{\frac{d}{2}}\},$ (74)
@@
Eigenvalues of $G$ when $d$ is odd.—In this case, for $i\in\{1,\cdots,(d-1)/2\}$, the nondegenerate eigenvalues are
@@
Eigenvalues of $G$ when $d$ is even.—In this case, the eigenvalues are the same as above, here for $i\in\{1,\cdots,d/2-1\}$, along with the two new eigenvalues $1+(L-1)|N_{d/2,d/2}|$ and $1-|N_{d/2,d/2}|$, which have multiplicities $1$ and L1L-1, respectively.
Now, let $\lambda_{\max}(G)$ be the maximum eigenvalue of the Gram matrix $G$. For both cases of even and odd $d$, we find
$\lambda_{\max}(G)=1+(L-1)\max_{i}|N_{i,d-i}|.$ (79)
Appendix C Parametrization of $q=(q_{0},\cdots,q_{d-1})$ for MEBs
For any dimension $d$, the nonzero entries of $N$ occurs when $k\oplus k^{\prime}=d$, as follows
$N_{kk^{\prime}}$ $=$ $\sum_{m=0}^{d-1}\omega^{k^{\prime}m}q_{m}=\sum_{m=0}^{d-1}\omega^{-km}q_{m}$
$=$ $\delta_{k}^{\textrm{R}}+i\delta_{k}^{\textrm{I}},$
where $\delta_{k}^{\textrm{R}}$ and $\delta_{k}^{\textrm{I}}$ are defined as the real and imaginary parts of $N_{kk^{\prime}}$, respectively, as
$\delta_{k}^{\textrm{R}}$ $=$ $\sum_{m=0}^{d-1}\cos\bigg{[}\frac{2\pi km}{d}\bigg{]}q_{m},$ (81)
$\delta_{k}^{\textrm{I}}$ $=$ $-\sum_{m=0}^{d-1}\sin\bigg{[}\frac{2\pi km}{d}\bigg{]}q_{m}.$ (82)
Accordingly, for the absolute values we have
$|N_{kk^{\prime}}|^{2}$ $=$ $[\delta_{k}^{\textrm{R}}]^{2}+[\delta_{k}^{\textrm{I}}]^{2}$
$=$ $\mu+2\sum_{m<m^{\prime}}^{d-1}\cos\bigg{[}\frac{2\pi k(m-m^{\prime})}{d}\bigg{]}q_{m}q_{m^{\prime}}.$
To illustrate the nature of this parametrization, we explicitly consider the cases $d=2,3$, and $4$, as follows.
d=2d=2.—In this case the only nonzero entry of $N$ is given by $N_{11}=\delta$, where $\delta=q_{0}-q_{1}$, such that $1\leq\delta\leq 1$. It follows therefore that
$q_{0}$ $=$ $(1+\delta)/2,$ (84)
$q_{1}$ $=$ $(1-\delta)/2.$
d=3d=3.—For $d=3$, the nonzero entries of $N$ are given by $N_{12}=N_{21}^{\ast}=\delta_{1}+i\delta_{2}$. Two independent parameters are given $\delta_{1}=(3q_{0}-1)/2$ and $\delta_{2}=\sqrt{3}(q_{2}-q_{1})/2$ where we have used them in Sec. II to plot the feasible region in $\delta_{2}\delta_{1}$-plane for $d=3$. Here
$-1/2\leq$ $\delta_{1}$ $\leq 1,$ (85)
$(-1+\delta_{1})/\sqrt{3}\leq$ $\delta_{2}$ $\leq(1-\delta_{1})/\sqrt{3}.$ (86)
We therefore have
$q_{0}$ $=$ $1/3+2\delta_{1}/3,$ (87)
$q_{1}$ $=$ $1/3-\delta_{1}/3-\delta_{2}/\sqrt{3},$
$q_{2}$ $=$ $1/3-\delta_{1}/3+\delta_{2}/\sqrt{3}.$
In this parametrization, one can write
$|N_{1,2}|=\sqrt{\delta_{1}^{2}+\delta_{2}^{2}}=\sqrt{(3\mu-1)/2}.$ (88)
d=4d=4.—When $d=4$, the nonzero entries of matrix $N$ are given by $N_{13}=N_{31}^{\ast}=\delta_{1}+i\delta_{2}$ and $N_{22}=\delta_{3}$. Therefore, three independent parameters are given by $\delta_{1}=q_{0}-q_{2}$, $\delta_{2}=q_{3}-q_{1}$, and $\delta_{3}=q_{0}-q_{1}+q_{2}-q_{3}$, such that
$-(1+\delta_{3})/2\leq$ $\delta_{1}$ $\leq(1+\delta_{3})/2,$ (89)
$-(1-\delta_{3})/2\leq$ $\delta_{2}$ $\leq(1-\delta_{3})/2,$ (90)
$-1\leq$ $\delta_{3}$ $\leq 1.$ (91)
It follows therefore that
$q_{0}$ $=$ $1/4+\delta_{1}/2+\delta_{3}/4,$
$q_{1}$ $=$ $1/4-\delta_{2}/2-\delta_{3}/4,$
$q_{2}$ $=$ $1/4-\delta_{1}/2+\delta_{3}/4,$
$q_{3}$ $=$ $1/4+\delta_{2}/2-\delta_{3}/4.$

### v2

@@ acknowledgment
Appendix A Derivation of the eigenvalues of $G$
@@
$D_{L}$ $=$ ${\mathrm{diag}}\{L-1,-1,\ldots,-1\},$ (77)
$\Lambda_{d-1}$ $=$ ${\mathrm{diag}}\{\lambda_{1},-\lambda_{1},\ldots,\lambda_{\lfloor\frac{d-1}{2}\rfloor},-\lambda_{\lfloor\frac{d-1}{2}\rfloor},\lambda_{\frac{d}{2}}\},$ (78)
@@
Eigenvalues of $G$ when $d$ is odd.—In this case, for $i\in\{1,\ldots,(d-1)/2\}$, the nondegenerate eigenvalues are
@@
Eigenvalues of $G$ when $d$ is even.—In this case, the eigenvalues are the same as above, here for $i\in\{1,\ldots,d/2-1\}$, along with the two new eigenvalues $1+(L-1)|N_{d/2,d/2}|$ and $1-|N_{d/2,d/2}|$, which have multiplicities $1$ and L1L-1, respectively.
For both cases of even and odd $d$, and $L\geq 2$, we have
$\lambda_{\max}(G)$ $=$ $1+(L-1)\max_{i}|N_{i,d-i}|,$ (83)
$\lambda_{\min}(G)$ $=$ $1-(L-1)\max_{i}|N_{i,d-i}|,$ (84)
for the largest and smallest eigenvalue of the Gram matrix $G$, respectively. The positive definiteness of the Gram matrix requires that $\max_{i}|N_{i,d-i}|<1/(L-1)$.